OpenStudy (anonymous):

ik this is the algebra not geomatry section but at this point im desprate dose anyone know how do do this... if so ill give medals. Solve the equation sin (x° - 20°) = cos 42° for x, where 0 < x < 90. A. 62 B. 22 C. 68 D. 28

OpenStudy (anonymous):

A

OpenStudy (nurali):

cos(42)=sin(90-42)=sin(48) sin(x-20)=sin(48) x-20=48 x=48+20 x=68 degrees

OpenStudy (anonymous):

thank u

OpenStudy (nurali):

Anytime.